Which of these reactions demonstrate the enthalpy of atom combination?

N(g)+3H(g)⟶NH3(g)N(g)+3H(g)⟶NH3(g)

C(graphite)+2O2(g)⟶CO2(g)C(graphite)+2O2(g)⟶CO2(g)

NO2(g)⟶N(g)+2O(g)NO2(g)⟶N(g)+2O(g)

2H2(g)+O2(g)⟶2H2O(g)2H2(g)+O2(g)⟶2H2O(g)

C(g)+O(g)⟶CO(g)
Answer #1

Solution:

Enthalpy of atom combination is defined as the change in enthalpy when one mole of compound is formed by the combination of its constituent atoms in gaseous state.

Hence, two reactions are showing enthalpy of atom combination,

N(g) + 3H(g) = NH3(g)

C(g) + O(g) = CO (g)